OpenStudy (anonymous):

What is the solution of this linear system? 3x + 5y = -1 2x − 5y = 16

OpenStudy (anonymous):

.3x + 5y = -1 2x - 5y = 16 Add the two equations: 5x = 15 or x = 3 Substitute for x in the first equation: 3*3 + 5y = -1 15 + 5y = -1 5y = -10 so y = -10/5 = -2 The solution is (x, y) = (3, -2)

OpenStudy (anonymous):

ahh thank you sooo much!!!

OpenStudy (okra77):

add the two equations yielding 5x=15 so x =3 then substitute into either equation to solve for y when you substitute you get y= -2

OpenStudy (owlcoffee):

We have the linear system: 3x + 5y = -1 2x − 5y = 16 Okay, let's see if we can simplify any of the x or y terms. an we can, see, on the first ecuation we have a +5y and on the second we have -5y. If we sum up the ecuations, we'll en up having: 3x = -1 2x = 16 = 5x=15 Now, let's find X. 5x=15 x=15/5 x=3 Now that we have one of the values, let's take any of the two initial ecuations and replace it: 3x + 5y = -1 ; x=3 => 3(3)-5y=-1 5y=-10 y=-10/5 y=-2. And theree we have it. We have found the solution for the linear system.

OpenStudy (anonymous):

thank you!!!!!!

OpenStudy (anonymous):

You are very welcome! If you need any more help let me know , I will try my best to help you (:

OpenStudy (anonymous):

i will for sure !thanks so much!

OpenStudy (anonymous):

My pleasure.

OpenStudy (anonymous):

*online

OpenStudy (anonymous):

Post them on openstudy sweety. I suggest you delete your number. Don't post your personal information online! Close this question and keep posting questions you need help with.

OpenStudy (anonymous):

okay